OpenStudy (aaronandyson):

the greater and least resultant of the two force are 7N & 3N respectively .'If each of the force is increased by 3 N and applied at 60 degrees.' Calculate the magnitude of the resultant force.

OpenStudy (aaronandyson):

@hartnn @inkyvoyd

OpenStudy (abhisar):

Ok, let the two forces be A and B. Now, the greatest resultant will be when the two forces are acting in same direction and least resultant will be when they are acting in opp direction.So you get two equations. A+B = 7 A-B = 3

OpenStudy (abhisar):

Solve the two equations to get the value of A and B. Now you can do the next part of the question.

OpenStudy (aaronandyson):

a = 10?

OpenStudy (abhisar):

2A = 10 So, A = 5

OpenStudy (aaronandyson):

b =2

OpenStudy (abhisar):

Yes, now do the next part of the question

OpenStudy (aaronandyson):

so each would be 8 and 5 righT? adding 3 to each?

OpenStudy (abhisar):

Yes.

OpenStudy (aaronandyson):

what next?

OpenStudy (abhisar):

Use the following formula to find their resultant \(\sf |\vec A + \vec B| = A^2+B^2 + 2AB \times cos \theta\)

OpenStudy (aaronandyson):

64 +25 right?

OpenStudy (abhisar):

Sorry, the formula is \(\sf \sqrt{A^2+B^2 + 2ABcos\theta}\)

OpenStudy (abhisar):

A=8 B=5 \(\sf \theta = 60\)

OpenStudy (aaronandyson):

whats the value of cos 60?

OpenStudy (abhisar):

0.5

OpenStudy (aaronandyson):

sqrt129

OpenStudy (abhisar):

Correct

OpenStudy (aaronandyson):

I need help with a few more questions n thanks.